OpenStudy (1645323):

Simplify: (2 + 3i)(4 - 5i) Select one: a. 18 + 6i b. 23 + 2i c. 6 - 2i d. 8 - 15i

OpenStudy (hoblos):

 (a+b)(c+d) = ac+ad+bc+bd so here (2 + 3i)(4 - 5i) = 2(4) + 2(-5i) + 3i(4) + 3i(-5i) = 8 -10i +12i - 15i2 where i2 = -1 = 8 + 2i +15 = 23 + 2i